# Clusterpath Gaussian Graphical Modeling

### v2

@@ 1 Introduction
Gaussian graphical models (GGM) are popular tools for summarizing conditional dependencies among $p$ variables. A GGM is a conditional dependency network where one refers to the variables as the nodes and the edges represent the conditional dependency relations among each pair of variables. Estimating GGMs is statistically challenging when the number of parameters ($p(p+1)/2$) is large relative to the sample size ($n$), leading to large estimation variability. Yet, such settings arise across many, diverse fields which has led to a flourishing area on regularized GGM estimation (e.g., Meinshausen & Bühlmann,, 2006; Peng et al.,, 2009; Cai et al.,, 2011). While much of the existing literature focuses on reducing the estimation variability through edge sparsity, our approach branches into a different direction by using node-clustering. We integrate convex clustering into the GGM framework to reduce estimation variability through estimation pooling for similar variables.
@@
Edge sparsity through, for instance, $\ell_{1}$-regularization has long formed the predominant choice to reduce dimensionality in GGMs (e.g., Yuan & Lin,, 2007; Yuan,, 2008; Friedman et al.,, 2008; Rothman et al.,, 2008). The idea is to sparsely estimate the precision matrix (i.e., the inverse of the covariance matrix)—the primary object of interest in GGMs—since conditional independencies between variable-pairs can be directly obtained from its sparsity pattern, or equivalently the absence of edges in the estimated GGM, offering an interpretability advantage. Several recent studies, however, point to important drawbacks when solely relying on edge sparsity as simplifying structure; namely weak detection capabilities in large-scale networks (Eisenach et al.,, 2020), interpretability issues for densely estimated GGMs with many variables (Grechkin et al.,, 2015), or inability to capture real world network structures such as hub nodes (Tarzanagh & Michailidis,, 2018) or block-structured graphs (Colombi et al.,, 2024).
@@
When GGMs face such challenges, one is often not interested in estimating conditional dependencies among the many observed variables but instead among a smaller number of (unobserved) clusters (also called communities) of original variables that share the same behavior. For instance, biologists estimate large-scale gene regulatory networks and cluster genes into pathways to unravel dependencies among them (e.g., Mestres et al.,, 2018; Shan et al.,, 2020), neuroscientists analyzing fMRI data routinely cluster voxels into regions of interest to learn interaction networks of brain activation (e.g., Pircalabelu & Claeskens,, 2020), financial analysts cluster company stocks into industry sectors to study how shocks can spread over the market by contagion (e.g., Wilms & Bien,, 2022). Cluster analysis is one of the most popular unsupervised learning methods to discover the underlying group structures in data. Variable clustering in GGMs not only offers simple, interpretable dependency networks but may also boost the dependency signals (Eisenach et al.,, 2020).
@@
To estimate GGMs with clustered variables, a recent yet growing interest arose in node-based dimensionality reduction. Initial proposals assume the clusters to be known a priori and incorporate this information in a regularization framework to encourage within-cluster over cross-cluster dependencies (e.g., Grechkin et al.,, 2015; Millington & Niranjan,, 2019). Domain-knowledge may, however, not always be available to impose a grouping, thereby still calling for unsupervised clustering procedures. Subsequent works learn the node clustering by decoupling the clustering task from the estimation of the GGM (e.g., Ambroise et al.,, 2009; Tan et al.,, 2015; Eisenach et al.,, 2020; Brownlees et al.,, 2022; Shi et al.,, 2024). Such a two-step procedure may however lead to suboptimal results thereby giving rise to a third stream that considers both tasks jointly (e.g., Pircalabelu & Claeskens,, 2020; Wilms & Bien,, 2022). Yet, to the best of our knowledge, the potential to leverage penalty structures popular in the literature on convex clustering (e.g., Pelckmans et al.,, 2005; Hocking et al.,, 2011; Lindsten et al.,, 2011; Chi et al.,, 2017 or, Weylandt et al.,, 2020; Chakraborty & Xu,, 2023 for more recent advances) to combine node clustering jointly with the estimation of the GGM is left largely unexplored, a notable exception being Yao & Allen, (2019).
@@
Our proposal is most closely related to Yao & Allen, (2019), Pircalabelu & Claeskens, (2020), and Wilms & Bien, (2022). There are two key distinctions with Yao & Allen, (2019): The first lies in the distance metric of the aggregation penalty: their approach does not account for the diagonal elements of the precision matrix, allowing the diagonal elements of clustered variables to differ. It is due to this property that the variable clustering is not retained when taking the inverse of the estimated precision matrix. Second, we show how the aggregation penalty of the CGGM estimator can be easily combined with other popular convex penalties such as a sparsity penalty. Furthermore, the procedure of Pircalabelu & Claeskens, (2020) results in a precision matrix with a blockdiagonal structure rather than a full block structure, whereas the regularizer of Wilms & Bien, (2022) requires side-information on the similarity of variables to guide node clustering.
@@ 2.1 Clustered GGMs
Our goal is to estimate the precision matrix $\mathbf{\Theta}$ (and hence the graph structure of the GGM) by encouraging clustering of the $p$ nodes in the graph. The $K$ new cluster variables $\xi_{1},\ldots,\xi_{K}$ then represent the average of the variables that belong to that cluster, with $K$ typically much smaller than $p$ to achieve dimensionality reduction in the parameters defining $\mathbf{\Theta}$. The edge weights among the clustered variables represent the conditional dependencies among the $K$ new cluster variables. Clustering of the variables in the graph corresponds to a block structure in the rows and columns of $\mathbf{\Theta}$, see the so-called G-block format introduced in Bunea et al., (2020) and discussed by Wilms & Bien, (2022) in the context of GGMs. In particular, for a partition $\{G_{1},\ldots,G_{K}\}$ of the variables $\{1,\ldots,p\}$ and corresponding $p\times K$ cluster membership matrix $\bf{U}$ with $u_{jk}=1$ if variable $j$ belongs to cluster $k$ and zero otherwise, there exists a $K\times K$ symmetric matrix $\mathbf{R}=(r_{k\ell})_{1\leq k,\ell\leq K}$, and a $p\times p$ diagonal matrix ${\bf A}=\text{diag}(a_{11}\mathbf{I},\ldots,a_{KK}\mathbf{I})$ such that the precision matrix can be written in $G$-block format given by
@@
A fundamental choice made throughout this paper is that $\mathbf{\Theta}$ should be positive definite. This holds if $\mathbf{R}+(\mathbf{U}^{\top}\mathbf{U})^{-1}\mathbf{A}$ is positive definite and $a_{kk}>0$, for more details, see Appendix A. As a consequence of this choice, the range of partial correlations that can be modeled may be limited in certain cases. For example, when $K=1$ and $\mathbf{\Theta}=r_{11}{\bf 1}{\bf 1}^{\top}+(1-r_{11}){\bf I}$, then $r_{11}$ is limited to the range $-1/(p-1)<r_{11}<1$ when requiring $\mathbf{\Theta}$ to be positive definite. Explicit ranges of allowed partial correlations are, however, difficult to formulate more generally for $K>1$ since this depends on the cluster structure.
In contrast to Bunea et al., (2020); Wilms & Bien, (2022), the block structure is not only reflected in the first part of decomposition (2.1) but also in the diagonal matrix ${\bf A}$. This means that we impose equal conditional variances of cluster members. An important implication of such an assumed block structure is that the clustering structure of $\mathbf{\Theta}$ is thereby retained when taking the inverse (e.g., Gower & Groenen,, 1991; Archakov & Hansen,, 2024). The approaches proposed by Yao & Allen, (2019), Pircalabelu & Claeskens, (2020), and Wilms & Bien, (2022) do not display this property. Section 4 explores the estimation of block-structured covariance matrices using CGGM through a numerical experiment.
@@
Finally, the block structure of the precision matrix in equation (2.1) may display sparsity, where some of the entries in $\mathbf{\Theta}$ are equal to zero. In particular, the absence of an edge in the clustered GGM between two cluster variables, for instance cluster variables one and two (i.e. $r_{12}=0$), is equivalent to cluster variables one and two being conditionally independent given all other variables, see Proposition 1 in Wilms & Bien, (2022). In turn, all members of cluster one are then conditionally independent of all members in cluster two.
@@ 2.2 Clusterpath Estimator
[…]
$\mathcal{P}(\mathbf{\Theta})=\lambda_{c}\sum_{j=1}^{p}\sum_{j^{\prime}=1}^{j-1}w_{jj^{\prime}}d_{jj^{\prime}}(\mathbf{\Theta})+\lambda_{s}\sum_{\begin{subarray}{c}j,j^{\prime}\\ j^{\prime}\neq j\end{subarray}}z_{jj^{\prime}}|\theta_{jj^{\prime}}|,$ (4)
@@
The first part in (4) represents the aggregation penalty penalizing differences $d_{jj^{\prime}}(\mathbf{\Theta})$ between columns $\bm{\theta}_{j}$ and $\bm{\theta}_{j^{\prime}}$ of the precision matrix $\mathbf{\Theta}$. Its role is thus to encourage estimation of a precision matrix with a $G$-block structure; or, put alternatively, a GGM with clustered variables. The second part represents the sparsity penalty on the unique off-diagonal elements of the precision matrix. Its role is to encourage estimation of precision matrices with zero off-diagonal entries; or, put alternatively, GGMs with edge-sparsity. The tuning parameters $\lambda_{c}$ and $\lambda_{s}$ control the degree of aggregation and sparsity respectively. The objective function is convex since each of the terms in (3) is convex in $\mathbf{\Theta}$ and $\mathbf{\Theta}$ lies in the convex cone of positive definite matrices.
@@
Focusing on the aggregation penalty, the weights $w_{jj^{\prime}}$, for every unique pair $1\leq j,j^{\prime}\leq p$, determine the fundamental attraction between two variables and are specified beforehand based on, for example, domain knowledge or the pairwise distances based on the sample precision matrix $\mathbf{S}^{-1}$. We further discuss the choice of weights in Section 2.3. If $d_{jj^{\prime}}(\bm{\Theta})=0$ and $\lambda_{c}$ is positive, then two vectors ${\bm{\theta}}_{j}$ and ${\bm{\theta}}_{j^{\prime}}$ have exactly the same elements while ignoring ${\theta}_{jj^{\prime}}$ and ${\theta}_{j^{\prime}j}$. The distance $d_{jj^{\prime}}(\mathbf{\Theta})$ disregards the difference between $\theta_{jj^{\prime}}$ and $\theta_{j^{\prime}j}$ as these are identical due to the symmetry of $\mathbf{\Theta}$. The penalty term $d_{jj^{\prime}}(\bm{\Theta})$ can thus be interpreted as a group lasso penalty (Yuan & Lin,, 2006), where each pair of variables $j<j^{\prime}$ forms a group whose elements are the differences between the corresponding entries in the columns $\bm{\theta}_{j}$ and $\bm{\theta}_{j^{\prime}}$ of the precision matrix. If all respective differences are put to zero, the estimated entries are identical, which in turn effectively blocks columns $j$ and $j^{\prime}$ in the precision matrix, or equivalently clusters nodes $j$ and $j^{\prime}$ in the GGM. As $\lambda_{c}$ increases (and $\lambda_{s}$ is fixed), estimated GGMs are obtained in which more and more variables are clustered, thereby resulting in a clusterpath from the $p$ original nodes (no clustering) until one clustered node (full clustering) in the GGM.
@@
The notion of applying a distance-based penalty to pairs of objects stems from convex clustering (Pelckmans et al.,, 2005; Hocking et al.,, 2011; Lindsten et al.,, 2011). In convex clustering, a copy of the data matrix is estimated while penalizing the distances between rows, thereby facilitating the clustering of observations. Building upon this framework, convex biclustering also clusters variables by adding a penalty on the distances between columns (Chi et al.,, 2017). Leveraging this notion for clustering variables via the precision matrix offers insights into the underlying conditional dependencies within the data.

### v3

@@ 1 Introduction
Gaussian graphical models (GGM) are popular tools for summarizing conditional dependencies among $p$ variables. A GGM is a conditional dependency network where one refers to the variables as the nodes and the edges represent the conditional dependency relations among each pair of variables. Estimating GGMs is statistically challenging when the number of parameters ($p(p+1)/2$) is large relative to the sample size ($n$), leading to large estimation variability. Yet, such settings arise across many, diverse fields which has led to a flourishing area on regularized GGM estimation (e.g., \@BBOPcitep\@BAP\@BBNMeinshausen and Bühlmann, 2006; Peng et al., 2009; Cai et al., 2011\@BBCP). While much of the existing literature focuses on reducing the estimation variability through edge sparsity, our approach branches into a different direction by using node-clustering. We integrate convex clustering into the GGM framework to reduce estimation variability through estimation pooling for similar variables.
@@
Edge sparsity through, for instance, $\ell_{1}$-regularization has long formed the predominant choice to reduce dimensionality in GGMs (e.g., \@BBOPcitep\@BAP\@BBNYuan and Lin, 2007; Yuan, 2008; Friedman et al., 2008; Rothman et al., 2008\@BBCP). The idea is to sparsely estimate the precision matrix (i.e., the inverse of the covariance matrix)—the primary object of interest in GGMs—since conditional independencies between variable-pairs can be directly obtained from its sparsity pattern, or equivalently the absence of edges in the estimated GGM, offering an interpretability advantage. Several recent studies, however, point to important drawbacks when solely relying on edge sparsity as simplifying structure; namely weak detection capabilities in large-scale networks \@BBOPcitep\@BAP\@BBN(Eisenach et al., 2020)\@BBCP, interpretability issues for densely estimated GGMs with many variables \@BBOPcitep\@BAP\@BBN(Grechkin et al., 2015)\@BBCP, or inability to capture real world network structures such as hub nodes \@BBOPcitep\@BAP\@BBN(Tarzanagh and Michailidis, 2018)\@BBCP or block-structured graphs \@BBOPcitep\@BAP\@BBN(Colombi et al., 2024)\@BBCP.
@@
When GGMs face such challenges, one is often not interested in estimating conditional dependencies among the many observed variables but instead among a smaller number of (unobserved) clusters (also called communities) of original variables that share the same behavior. For instance, biologists estimate large-scale gene regulatory networks and cluster genes into pathways to unravel dependencies among them (e.g., \@BBOPcitep\@BAP\@BBNMestres et al., 2018; Shan et al., 2020\@BBCP), neuroscientists analyzing fMRI data routinely cluster voxels into regions of interest to learn interaction networks of brain activation (e.g., \@BBOPcitep\@BAP\@BBNPircalabelu and Claeskens, 2020\@BBCP), financial analysts cluster company stocks into industry sectors to study how shocks can spread over the market by contagion (e.g., \@BBOPcitep\@BAP\@BBNWilms and Bien, 2022\@BBCP). Variable clustering in GGMs not only offers simple, interpretable dependency networks but may also boost the dependency signals \@BBOPcitep\@BAP\@BBN(Eisenach et al., 2020)\@BBCP.
@@
To estimate GGMs with clustered variables, a recent yet growing interest arose in node-based dimensionality reduction. While edge sparsity simplifies the graph structure by removing (noisy) edges, variable clustering achieves a simpler graph by aggregating similar nodes. Note that reducing the number of nodes through aggregation automatically reduces the number of edges in the graph as well. Initial proposals assume the clusters to be known a priori and incorporate this information in a regularization framework \@BBOPcitep\@BAP\@BBN(e.g., Grechkin et al., 2015; Millington and Niranjan, 2019)\@BBCP. Nevertheless, domain-knowledge may not always be available to impose a grouping, calling for unsupervised clustering procedures. Subsequent works learn the node clustering by decoupling the clustering task from the estimation of the GGM (e.g., \@BBOPcitep\@BAP\@BBNAmbroise et al., 2009; Tan et al., 2015; Eisenach et al., 2020; Brownlees et al., 2022; Shi et al., 2024\@BBCP). Such a two-step procedure may however lead to suboptimal results, giving rise to a third stream that considers both tasks jointly, thereby successfully leveraging a block-structure in the precision matrix to achieve node clustering (e.g., \@BBOPcitep\@BAP\@BBNYao and Allen, 2019; Pircalabelu and Claeskens, 2020; Wilms and Bien, 2022\@BBCP). Yet, to the best of our knowledge, the potential to leverage penalty structures popular in the literature on convex clustering (e.g., \@BBOPcitep\@BAP\@BBNPelckmans et al., 2005; Hocking et al., 2011; Lindsten et al., 2011; Chi et al., 2017\@BBCP or, \@BBOPcitep\@BAP\@BBNWeylandt et al., 2020; Chakraborty and Xu, 2023\@BBCP for more recent advances) to combine node clustering jointly with the estimation of the GGM is left largely unexplored, with \@BBOPcite\@BAP\@BBNYao and Allen (2019)\@BBCP being the exception.
@@
Our proposal is most closely related to \@BBOPcite\@BAP\@BBNYao and Allen (2019)\@BBCP, \@BBOPcite\@BAP\@BBNPircalabelu and Claeskens (2020)\@BBCP, and \@BBOPcite\@BAP\@BBNWilms and Bien (2022)\@BBCP. There are two key distinctions with \@BBOPcite\@BAP\@BBNYao and Allen (2019)\@BBCP: The first lies in the distance metric of the aggregation penalty: their approach does not account for the diagonal elements of the precision matrix, allowing the diagonal elements of clustered variables to differ. It is due to this property that the variable clustering is not retained when taking the inverse of the estimated precision matrix. Second, we show how the aggregation penalty of the CGGM estimator can be easily combined with other popular convex penalties such as a sparsity penalty. Furthermore, the procedure of \@BBOPcite\@BAP\@BBNPircalabelu and Claeskens (2020)\@BBCP results in a precision matrix with a blockdiagonal structure rather than a full block structure, whereas the regularizer of \@BBOPcite\@BAP\@BBNWilms and Bien (2022)\@BBCP requires side-information on the similarity of variables to guide node clustering.
@@ 2.1 Clustered GGMs
Our goal is to estimate the precision matrix $\mathbf{\Theta}$ (and hence the graph structure of the GGM) by encouraging clustering of the $p$ nodes in the graph. The $K$ new cluster variables $\xi_{1},\ldots,\xi_{K}$ then represent the average of the variables that belong to that cluster, with $K$ typically much smaller than $p$ to achieve dimensionality reduction in the parameters defining $\mathbf{\Theta}$. The edge weights among the clustered variables represent the conditional dependencies among the $K$ new cluster variables. Clustering of the variables in the graph corresponds to a block structure in the rows and columns of $\mathbf{\Theta}$, see the so-called G-block format introduced in \@BBOPcite\@BAP\@BBNBunea et al. (2020)\@BBCP and discussed by \@BBOPcite\@BAP\@BBNWilms and Bien (2022)\@BBCP in the context of GGMs. In particular, for a partition $\{G_{1},\ldots,G_{K}\}$ of the variables $\{1,\ldots,p\}$ and corresponding $p\times K$ cluster membership matrix $\bf{U}$ with $u_{jk}=1$ if variable $j$ belongs to cluster $k$ and zero otherwise, there exists a $K\times K$ symmetric matrix $\mathbf{R}=(r_{k\ell})_{1\leq k,\ell\leq K}$, and a $p\times p$ diagonal matrix ${\bf A}=\text{diag}(a_{11}\mathbf{I},\ldots,a_{KK}\mathbf{I})$ such that the precision matrix can be written in $G$-block format given by
@@
A fundamental choice made throughout this paper is that $\mathbf{\Theta}$ should be positive definite. This holds if $\mathbf{R}+(\mathbf{U}^{\top}\mathbf{U})^{-1}\mathbf{A}$ is positive definite and $a_{kk}>0$, for more details, see Appendix A. As a consequence of this choice, the range of partial correlations that can be modeled may be limited in certain cases. For example, when $K=1$ and $\mathbf{\Theta}=r_{11}{\bf 1}{\bf 1}^{\top}+(1-r_{11}){\bf I}$, then $r_{11}$ is limited to the range $1/(p-1)<r_{11}<1$ when requiring $\mathbf{\Theta}$ to be positive definite. Explicit ranges of allowed partial correlations are, however, difficult to formulate more generally for $K>1$ since this depends on the cluster structure.
In contrast to \@BBOPcite\@BAP\@BBNBunea et al. (2020); Wilms and Bien (2022)\@BBCP, the block structure is not only reflected in the first part of decomposition (2.1) but also in the diagonal matrix $\bf A$}. This means that we impose equal conditional variances of cluster members. An important implication of such an assumed block structure is that the clustering structure of $\mathbf{\Theta}$ is thereby retained when taking the inverse \@BBOPcitep\@BAP\@BBN(e.g., Gower and Groenen, 1991; Archakov and Hansen, 2024)\@BBCP. The approaches proposed by \@BBOPcite\@BAP\@BBNYao and Allen (2019)\@BBCP, \@BBOPcite\@BAP\@BBNPircalabelu and Claeskens (2020)\@BBCP, and \@BBOPcite\@BAP\@BBNWilms and Bien (2022)\@BBCP do not display this property. Section 4 explores the estimation of block-structured covariance matrices using CGGM through a numerical experiment.
@@
Finally, the block structure of the precision matrix in Equation (2.1) may display sparsity, where some of the entries in $\mathbf{\Theta}$ are equal to zero. In particular, the absence of an edge in the clustered GGM between two cluster variables, for instance cluster variables one and two (i.e. $r_{12}=0$), is equivalent to cluster variables one and two being conditionally independent given all other variables, see Proposition 1 in \@BBOPcite\@BAP\@BBNWilms and Bien (2022)\@BBCP. In turn, all members of cluster one are then conditionally independent of all members in cluster two.
@@ 2.2 Clusterpath Estimator
where $\log|\cdot|$ denotes the logarithm of the determinant, $\operatorname*{tr}(\cdot)$ is the trace, $\cdot\succ 0$ denotes a positive definite matrix, and $\mathcal{P}(\mathbf{\Theta})$ is the penalty term. Throughout the paper we take
$\mathcal{P}(\mathbf{\Theta})=\lambda_{c}\sum_{j=2}^{p}\sum_{j^{\prime}=1}^{j-1}w_{jj^{\prime}}d_{jj^{\prime}}(\mathbf{\Theta})+\lambda_{s}\sum_{\begin{subarray}{c}j,j^{\prime}\\ j^{\prime}\neq j\end{subarray}}z_{jj^{\prime}}|\theta_{jj^{\prime}}|,$ (4)
@@
The first part in (4) represents the aggregation penalty penalizing differences $d_{jj^{\prime}}(\mathbf{\Theta})$ between columns $\bm{\theta}_{j}=(\theta_{1j},...,\theta_{pj})^{\top}$ and $\bm{\theta}_{j^{\prime}}$ of the precision matrix $\mathbf{\Theta}$. Its role is thus to encourage estimation of a precision matrix with a $G$-block structure; or, put alternatively, a GGM with clustered variables. The second part represents the sparsity penalty on the unique off-diagonal elements of the precision matrix with sparsity weights $z_{jj^{\prime}}$, defined in Appendix A.1. Its role is to encourage estimation of precision matrices with zero off-diagonal entries; or, put alternatively, GGMs with edge-sparsity. The tuning parameters $\lambda_{c}$ and $\lambda_{s}$ control the degree of aggregation and sparsity respectively. The objective function is convex since each of the terms in (3) is convex in $\mathbf{\Theta}$ and $\mathbf{\Theta}$ lies in the convex cone of positive definite matrices.
@@
Focusing on the aggregation penalty, the weights $w_{jj^{\prime}}$, for every unique pair $1\leq j,j^{\prime}\leq p$, determine the fundamental attraction between two variables and are specified beforehand based on, for example, domain knowledge or the pairwise distances based on the sample precision matrix $\mathbf{S}^{-1}$. We further discuss the choice of weights in Section 2.3. If $d_{jj^{\prime}}(\bm{\Theta})=0$ and $\lambda_{c}$ is positive, then two vectors ${\bm{\theta}}_{j}$ and ${\bm{\theta}}_{j^{\prime}}$ have exactly the same elements while ignoring ${\theta}_{jj^{\prime}}$ and ${\theta}_{j^{\prime}j}$. The distance $d_{jj^{\prime}}(\mathbf{\Theta})$ disregards the difference between $\theta_{jj^{\prime}}$ and $\theta_{j^{\prime}j}$ as these are identical due to the symmetry of $\mathbf{\Theta}$. We do also penalize the differences between entries on the diagonal, since the block structure must also be reflected in the diagonal of the precision matrix, see Equation (2.1). The penalty term $d_{jj^{\prime}}(\bm{\Theta})$ can thus be interpreted as a group lasso penalty \@BBOPcitep\@BAP\@BBN(Yuan and Lin, 2006)\@BBCP, where each pair of variables $j<j^{\prime}$ forms a group whose elements are the differences between the corresponding entries in the columns $\bm{\theta}_{j}$ and $\bm{\theta}_{j^{\prime}}$ of the precision matrix. If all respective differences are put to zero, the estimated entries are identical, which in turn effectively blocks columns $j$ and $j^{\prime}$ in the precision matrix, or equivalently clusters nodes $j$ and $j^{\prime}$ in the GGM. The shrinkage of the distances between the clusters pushes the corresponding entries in the precision matrix closer to each other across clusters, thereby introducing bias; we accommodate this through a common ‘post-selection’ refitting step in the regularization literature \@BBOPcitep\@BAP\@BBN(Belloni and Chernozhukov, 2013)\@BBCP, see Section 2.3.4. As $\lambda_{c}$ increases (and $\lambda_{s}$ is fixed), estimated GGMs are obtained in which more and more variables are clustered, thereby resulting in a clusterpath from the $p$ original nodes (no clustering) until one clustered node (full clustering) in the GGM.
@@
The notion of applying a distance-based penalty to pairs of objects stems from convex clustering \@BBOPcitep\@BAP\@BBN(Pelckmans et al., 2005; Hocking et al., 2011; Lindsten et al., 2011)\@BBCP. In convex clustering, a copy of the data matrix is estimated while penalizing the distances between rows, thereby facilitating the clustering of observations. Building upon this framework, convex biclustering also clusters variables by adding a penalty on the distances between columns \@BBOPcitep\@BAP\@BBN(Chi et al., 2017)\@BBCP. Leveraging this notion for clustering variables via the precision matrix offers insights into the underlying conditional dependencies within the data.